√(a-1)−5=1 solve equation and check for extraneous solution
Answer by dabanfield(803)

√(a-1)−5=1 solve equation and check for extraneous solution
sqrt(a-1) - 5 = 1
sqrt(a-1) = 1 + 5
sqrt(a-1) = 6
Square both sides:
a-1 = 36
a = 37
37 is not extraneous.